B is inversely proportional to a. If b=9 when a=3, find b when a=3/2
Answer by stanbon(75887)

B is inversely proportional to a. If b=9 when a=3, find b when a=3/2
-----
B = k/a
----
Solve for k using "B = 9 when a = 3"::
9 = k/3
k = 27
-----
Equation:
B = 27/a
----
Now, find B when a = 3/2::
B = 27/(3/2)
B = 18
------------
Cheers,
Stan H.
---------